OpenStudy (anonymous):

Write an equation in slope-intercept form of the line with the given slope and point: slope = 3 and (1, -2). A. y = 3x - 5 B. y = 3x - 2 C. y = 3x + 1 D. y = 3x + 3

OpenStudy (alexandervonhumboldt2):

y=3x-5

OpenStudy (anonymous):

im not doubting you but do you have an equation

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

y=3x-5 is the equation

OpenStudy (anonymous):

lol i mean how did you find out the equation

OpenStudy (alexandervonhumboldt2):

ok so the slope is 3 so y=3x+b substitute (1, -2) -2=3+b b=-5 y=3x-5

OpenStudy (anonymous):

ok i was just making sure you didnt just write a random answer

OpenStudy (anonymous):

that is good that you asked for an explanation. It shows that you are trying to learn.

OpenStudy (alexandervonhumboldt2):

yeah thx

OpenStudy (anonymous):

thanks guys